Reports 1-1 of 1 Clear search Modify search
VIS (General)
satoru.ikeda - 16:33 Thursday 19 August 2021 (17964) Print this report
Installation of temperature control equipment

Temperature control equipment was brought into the machine room tent in the KAGRA mine.

The installation process will be done later.

Non-image files attached to this report
Comments to this report:
satoru.ikeda - 16:10 Friday 20 August 2021 (17972) Print this report

[Ikeda, Yasui-san]

We have installed temperature control devices in the VIS miniracks of PRM,PR3,PR2,BS,SR2,SR3,SRM.
The cables are not connected yet. We will install the cables next time.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 15:53 Monday 23 August 2021 (17982) Print this report

Installed temperature control devices in IXV and IYV.
Installed in VIS mini rack and connected power supply 24V and LAN cable.
Added the MEDM screen and confirmed that the temperatures of ETMX and ETMY are being acquired.
The heater will be installed in the next corrugated tube.

MEDM:
 [sitemap]-[VIS]-[VIS_TEMPERATURE]

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:02 Tuesday 24 August 2021 (17996) Print this report

[Ikeda, Yasui-san, Takahashi-san]

We have installed the ribbon heater on ITMX.
However, it is not completed yet.
There were two problems.
I don't have enough space for a 100V AC power supply, and when I do, the cable length is not long enough.
 => The cable length is not long enough.
Instead of using a single 2m ribbon heater, we use a combination of 2m and 1m.
 Therefore, it is necessary to make the mouth of the table tap two-way.

 => I'm planning to buy a two-way outlet.

[Ikeda, Yasui-san, Fabian-san, Hirata-san]

Finished connecting cables for SRM, SR3, SR2, BS, PR2.
We have not yet installed the ribbon heater.

We found one problem.
The temperature of PT100 on SRM showed an error message.
=> I will check it next time.
The SRM and SR3 areas are using orange cables.
=> It needs to be changed to 24V. Measure the voltage and if it is 24V, label it.
  If the voltage is 18V, change it to 24V and put a label on it.
 

Photos: TemperatureControl

 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:09 Wednesday 25 August 2021 (18015) Print this report

[Ikeda, Fabian-san, Hirata-san, Yasui-san].

We connected the power and communication cables of the temperature control equipment and the thermistor (PT-100) to the PRM and PR3.
We checked from MEDM and confirmed that the communication is normal.

The temperature of PT100 was displayed as an error in SRM.
=> Checked and found that the problem is on the chamber side rather than the DSUB cable.
 To be confirmed next time.

The orange cable is used in SRM and SR3 area.
 The SRM and SR3 areas are using orange cables.

=> First, we labeled which devices the cables are connected to.
 They were supplied with 18V. 
 The devices connected to each chamber were two of the stepper motors and one of the OpLevs.
 These stepper motors needed to be powered by 24V.
 However, there are 24V table taps installed in the SRM, SR2, and SR3 Racks, but they are not supplied with power.
 Therefore, we need to have 24V power supplied to the table tap first.

[Ikeda, Yasui-san]

We have installed ribbon heaters (1m, 2m) to ITMX, but we have not connected the heaters yet due to some problems.

We don't have enough space for AC100V power supply, and when we do, the cable length is not long enough.
 => I ordered extension cables and such.

I found one problem.
The control signal from the KT4R to the power regulator is variable and constant at 20mA, and the heater is always on, so the heater is not controlled.
=> Plan to review parameters of KT4R.

Photo: TemperatureControl

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:03 Thursday 26 August 2021 (18025) Print this report

[Ikeda,Yasui-san]

We installed a temperature control device on the ETMX.
=> I haven't installed the heater yet.

The heater is not controlled because the control signal from KT4R to the power adjustment device is variable and constant 20mA and the heater is always ON.
=> We reviewed the parameters of KT4R and did not find any problem.
 Plan to create a test environment and verify again.
 

Photo: TemperatureControl

Images attached to this comment
Non-image files attached to this comment
fabian.arellano - 16:21 Monday 30 August 2021 (18072) Print this report

With Hirata-san.

On Thursday 26th of August we fixed the thermistor connection problem.

Pin 1 in the in-vacuum connector at the flange wasn't making appropriate contact. All the pins had to be removed and new ones were crimped. After this operation it worked well.

satoru.ikeda - 18:05 Monday 30 August 2021 (18076) Print this report

[Ikeda, Yasui-san]

Installed temperature control devices on ETMY.

Temperature control devices are now installed in PRM, PR2, PR3, BS, SR2, SR3, ITMX, ITMY, ETMX, and ETMY.
We will continue to install the heaters.

We checked if the ribbon heater can be installed around the PRM chamber, but the size is larger than expected (about 4.8m).
The equipment such as OpLev is bare, so we will consult with them a little before proceeding.

The control signal from KT4R to the power adjustment device is variable and constant at 20mA, so the heater is always in ON state.
=> The cause has been found and will be corrected next time.
 The cause was a setting error in the power regulator.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:11 Monday 30 August 2021 (18077) Print this report

[Ikeda, Hirata-san, Yasui-san]

We changed the power supply for the stepper motor drivers of SR2, SR3 and SRM from 18V to 24V.
BS had a 24V cable connected to it.

All temperature control devices are now connected to 24V.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:53 Tuesday 31 August 2021 (18086) Print this report

[Ikeda, Yasui-san]

We replaced the ribbon heater of ITMX with polyimide tape.
We also checked the operation of the ribbon heater by actually increasing the temperature.

With the temperature at 23.1°C when it was off, I set the heater to 23.1°C and checked if it was stable.
Immediately after the power was turned on, it showed a transient response, but the temperature stabilized, so we raised the temperature to 24°C and checked it.
We confirmed that the temperature was almost stable and could be controlled at 24℃±0.1℃.

However, after about 45 minutes, the power of the electrical regulator suddenly turned off.
I'm afraid to turn it on again, so I plan to consult with Takahashi-san and replace it with a new one next time to try.
Suspicion of initial failure?
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 12:35 Wednesday 01 September 2021 (18095) Print this report

[Ikeda, Yasui-san]

Yesterday, I investigated the case of ITMX's power regulator being turned off.
First, I asked Yoshimura-san to check the IXV wall outlets, breakers and other equipment.
=> No problems were found.

Next, we used the new power regulator to check its operation.
=> Unlike yesterday, the test was completed without any problems.

Based on the above, we believe that the problem is with the power regulator itself.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:13 Monday 06 September 2021 (18162) Print this report

[Ikeda, Yasui-san]

Preparing for installation of ribbon heater on ETMX

In the morning we prepared for the installation of the ribbon heater in ETMX.
We found out that we need to climb on top of the clean booth to pull the cable into the clean booth on the mezzanine floor.
To check the procedure, I only installed the power regulator in the morning and stopped the work in the afternoon.
We plan to work with Takahashi-san later in the day.

Fixing the screws on the DSUB9P connector of the thermistor (PT-100)

I checked the screw fastening again.
It seems that if you close all the screws on one side, the other side cannot be fixed.
I was able to fix it by alternately rotating one side one turn and the other one one turn.
I fixed the DSUBs of all thermistors in this way.
=> PRM,PR2,PR3,BS,SR2,SR3,SRM
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:07 Tuesday 07 September 2021 (18180) Print this report

[Ikeda, Takahashi-san, Yasui-san]

Installed a ribbon heater in ETMX.
We confirmed that the temperature rises from room temperature to 25.0℃.

The AC100V is currently unplugged.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:33 Thursday 09 September 2021 (18216) Print this report

Investigation of voltage and temperature to ribbon heaters

As a safety measure for the ribbon heater, we investigated the settings necessary to keep the upper temperature limit of the ribbon heater at about 40°C.

The upper limit of the output voltage can be changed by changing the output gradient setting volume (SLOPE) of the power regulator.
This is the maximum voltage (100% = maximum 100V) when 20mA (100% output) is given by the control line from KT4R.

The power regulator was set to MANUAL setting and the output voltage was measured by changing the SLOPE volume.
The test was spread out in a cabinet and checked.Although not shown in the photo, I placed an aluminum sheet to keep the heat in.
The operation will be checked in the chamber based on this data.

Voltage and Temperature
SLOPE Output voltage [V] Temperature [°C]
3 7.28 27.4
4 14.04 32.0
5 28.6 70 over


From the above, the output can be suppressed by setting the SLOP of the power regulator side to 4.
However, it is necessary to make a final decision depending on the heat dissipation situation at the time of actual installation.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:27 Monday 13 September 2021 (18241) Print this report

[Ikeda]

 In the morning, I got the ITMX ribbon heater ready for 24-hour use.
 I set the SLOP of the power regulator to 2 and set the target temperature to 40°C to see at what temperature the actual heater would reach steady state.
 After 5 hours of operation, the heater reached steady state at about 30.2℃.

 The target temperature was changed to 0°C and the work was finished.

[Ikeda, Takahashi-san, Yasui-san]

 In the afternoon, we wrapped the BS and SR2 ribbon heaters around the corrugated tube.
 Thermal insulation and thermocouple installation will be done tomorrow.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:01 Tuesday 14 September 2021 (18252) Print this report

[Ikeda, Yasui-san]

We have installed the ribbon heaters for SRM and SR3.
We also wired the BS, SR2, SR3 and SRM ribbon heaters and installed thermocouples in SR3.

We connected a power regulator to the ribbon heater of SR3 and tested to raise the temperature, but the temperature did not rise much in SLOP2.
Next time, we will install thermocouples and recheck SR3.
 

Images attached to this comment
Non-image files attached to this comment
shinji.miyoki - 10:19 Thursday 16 September 2021 (18276) Print this report

I guess some activities and sets of instruments in the EXV area seem to enhance the temp in EXV by 0.4C on 7th September. I guess we can reduce the temp by a stop of a small heater.

However, we need to know the optimum temp in EXV and other V areas.

satoru.ikeda - 16:01 Tuesday 21 September 2021 (18319) Print this report

[Ikeda, Yasui-san]

Installed ribbon heaters on BS, SR2, SR3 and SRM.
SR3 was a little off, so I had to reinstall it.
No heat retention with aluminum for now.

Continued to check the operation of the ribbon heater on SR3.
With SLOPE2 and 3 on the power regulator, the temperature hardly rose.
Next, we changed to SLOP4 and set the target temperature to 30°C. We checked the temperature for about 2 hours, but it only rose to 27.2°C.
Since this is lower than we had thought, we will review the SLOPE value and test again later.

The other BS, SR2 and SRM will be tested for temperature rise as well.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:17 Wednesday 22 September 2021 (18335) Print this report

The following work was done related to temperature control devices.

ITMX:
 I stored the AC connection connector of the ribbon heater in a case.
 It was put into a commercially available case, but the hole was too big and needed to be sealed.
 Will figure out how to stop it with waterproof tape.
 The remaining work for IX is to replace the above waterproofing measures with aluminum tape.

Continuing from yesterday, we tested the ribbon heater on SR3.
 First, we set the power regulator to SLOPE 5 and set the target temperature to 30℃.
 The temperature rose quickly and was about to exceed 30℃, so we changed the setting in the middle of the test and continued.
 The power regulator was set to SLOPE 4.5, and the target temperature was set to 30℃.
 After one hour, the temperature reached 29.9℃.
 There was no further increase in temperature, so I judged that the target was achieved.
 We will try Type-B at 4.5 next time.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:30 Tuesday 28 September 2021 (18375) Print this report

[Ikeda, Hirata-san]

Wrapped aluminum tape around the ribbon heaters of SR3 and SRM.
A shorter piece of tape was split and attached over the thermocouple to allow for maintenance.

We checked the temperature rise after wrapping the aluminum tape with SR3.
Against the target temperature of 40°C, the temperature rose to 30.7°C after about 2 hours.
This is higher temperature than the previous time without aluminum tape (29.9℃).

We plan to work on IX, BS, and SR2 at a later date.

[Remaining work on temperature control devices]

IY, EY: Installed ribbon heaters, installed thermocouples, rolled aluminum tape, installed power regulator, temperature rise test
BS, SR2, EX: Aluminum tape winding, installation of power regulator, temperature rise test
SRM: Installation of power regulator, temperature rise test
EX: Aluminum tape wrapping, temperature rise test
SR3: Place the power regulator on the VIS mini rack and the job is done!
IX: Wrapping aluminum tape
PR2, PR3, PRM: Installation if necessary

[Main work]
Installation of power regulator: Waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:07 Wednesday 29 September 2021 (18395) Print this report

[Ikeda, Yasui-san, Nakagaki-san]

We put aluminum tape on the ribbon heaters of BS, SR2 and ITMX.
Also, We did a temperature rise test for SRM.
We set the target temperature to 40℃ and checked the temperature change.
After 2 hours, the temperature rose to 28.3℃. The SLOPE of the temperature controller was 4.5, so the result was lower than SR3.
At the same time, I screwed up the stepper motor driver installed in SRM's VIS minirack, which is a different matter from the temperature control device.

[Remaining work on temperature control equipment]

IY, EY: Installed ribbon heaters, installed thermocouples, rolled aluminum tape, installed power regulator, temperature rise test
EX: Wrapping aluminum tape, temperature rise test
IX: Available
BS, SR2, SRM: Installation of power regulator, temperature rise test
SR3: Place power regulator on VIS mini-rack and job is done
PR2, PR3, PRM: Installation if necessary

[Main work].
Installation of power regulator: Waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:31 Monday 04 October 2021 (18439) Print this report

[Ikeda, Yasui-san]

Worked on installing the next temperature controller.

ETMX
 Wrapped aluminum tape around the ribbon heater and set it to be available for 24 hours.

 We set the SLOPE of the power regulator to 4 and the target temperature to 40°C to check the temperature rise.
 After checking for about 5 hours, the temperature rose to 34.0℃.

SR2
 The temperature did not rise even after changing SLOPE to 5.5.
  No current was flowing in the control line connected to the power regulator.
  The connection of the control line may have been broken inside the temperature control box.
  We will remove the temperature control from the VIS minirack next time to check the inside.

BS
 We set the SLOPE of the power regulator to 4.5 and the target temperature to 40°C to check the temperature rise.
 After checking for about 2 hours, the temperature rose to 32.0℃.

All the target temperatures were changed to 0°C, and the work was completed.

[Remaining work on temperature control devices]

IY, EY: Installation of ribbon heaters, installation of thermocouples, wrapping aluminum tape, installation of power regulator, temperature rise test
IX, EX: Ready for use
SR2: Temperature rise test in progress (there is a problem that the temperature does not rise)
BS, SR3, SRM: Temperature rise test completed, installed after delivery of power regulator (November)
PR2, PR3, PRM: Ribbon heaters have been purchased and will be installed after confirming the necessity of installation.

[Main work]
Installation of power regulator: Waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:35 Tuesday 05 October 2021 (18464) Print this report

[Ikeda, Hirata-san, Yano-san]

We have confirmed the problem with the temperature controller of SR2.

[Description of the problem]
The temperature of the ribbon heater does not rise even when the temperature is set.
 
[Description]
Opened the lid of the temperature controller (S2113165) and checked the cable connections.
=> No abnormality was found in the cable between the KT4R terminal and the power regulator.
In the previous check, the output of the control line to the power regulator was checked, but there was no output at all. The output voltage to the ribbon heater does not change.
The output voltage to the ribbon heater did not change, and the temperature was set correctly by connecting to the KT4R (S2112834) used as a thermometer in the chamber instead of the KT4R (S2113037) on the heater side.
From the above, I suspect that the KT4R may be initially defective.

[We will check next time.]
We are planning to replace it with a spare KT4R.

Non-image files attached to this comment
satoru.ikeda - 16:19 Wednesday 06 October 2021 (18478) Print this report

[Ikeda, Yano-san(Help for assembly)]

The problem with the temperature controller of SR2 has been fixed.

[Description of the problem]
The temperature of the ribbon heater did not rise even when the temperature was set.
 
[Description of the problem]
The temperature of the ribbon heater did not rise when the temperature was set.
 The removed KT4R (S2112836) was moved to the 2nd floor of Hokubu-Kaikan.
When I checked the cable connections again, I found that they were connected to the terminals but had been replaced.
 It turned out that it was not a KT4R problem, but a connection error.
 The terminal side is soldered and difficult to repair, so I rewired the KT4R side.
 I connected the - and + cables to #7:+ and #8:- respectively, which are connected to KT4R. (Normally, the + and - cables are connected as shown.)

[Test]
We continued the temperature rise test of SR2.

We set the target temperature to 40°C with SLOPE 4.5, and checked the temperature rise.
The temperature rose to 27.4°C after 2 hours.
We think we need to change to SLOPE 5.5, the same as SR2, because the temperature reached by SLOPE 4.5 is too low.

[Remaining work on temperature control equipment]

IY, EY: Installation of ribbon heater, installation of thermocouple, wrapping aluminum tape, installation of power regulator, temperature rise test
IX, EX: Available
BS, SR2, SR3, SRM: Temperature rise test completed, installation after delivery of power regulator (November)
PR2, PR3, PRM: Ribbon heaters already purchased, to be installed after confirming the necessity of installation

[Main work]
Installation of power regulator: Waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:06 Tuesday 26 October 2021 (18695) Print this report

[Ikeda, Nakagaki-san, Yasui-san, Ushiba-san(Remote)]

Installed ribbon heaters on PRM and PR2.
PRM is all installed.
PR2 could not be installed until the end because of the missing three-prong outlet, lack of parts, and lack of time.
The thermometer and two seals installed in the PR2 chamber were moved down a bit after consulting with Ushiba-san, Yokozawa-san and Kimura-san respectively.

Performed temperature rise test on PRM.
We set the SLOPE of the temperature controller to 4 and the target temperature to 40°C. We heated the unit for about an hour and a half, and the temperature rose to 34.1°C.
PR2 will be tested again later.

We checked the value of TM_WIT in PR3 before and after the work and found that the value fluctuated as shown in K-Log #18691, but it was due to LED lighting and was not affected by this work.

[Remaining work on temperature control equipment]

IY, EY: Installed ribbon heaters, installed thermocouples, wrapped aluminum tape, installed power regulator, temperature rise test
IX, EX, PRM: Available
BS, SR2, SR3, SRM: Temperature rise test completed, power regulator installed after delivery (November)
PR2: Need to re-purchase three-prong outlet, otherwise complete, temperature rise test not yet conducted
PR3: Ribbon heater purchased, to be installed after confirming need for installation.

[Main work]
Installation of power regulator: waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 18:09 Wednesday 27 October 2021 (18714) Print this report

[Ikeda, Yasui-san(pm)]

PR2 and SR3 can now be used.

[Work description]
Installed power regulators on BS, SR3, and SRM.
(The one installed in SR3 and SRM used the one in Mitaka. I cleaned them.)
We also installed a three-pronged outlet on PR2.


[Test results]
We tested the temperature rise of PR2 and SR3.
All of them were tested with SLOPE set to 4.5 and target temperature set to 40°C.
The temperature of PR2 rose to 35.5°C after 3 hours.
SR3 rose to 34.2°C after 1.5 hours.

Finally, the target temperature was changed to 0°C and the work was completed.

[Remaining work on temperature control equipment]

IY, EY, PR3: Installed ribbon heaters, installed thermocouples, wound aluminum tape, installed power regulator, temperature rise test
IX, EX, PRM, PR2, SR3: Available
BS, SR2, SRM: AC100V outlet not connected, final check to be conducted

[Main work contents]
Installation of power regulator: waiting for power regulator to be obtained
Temperature rise test: Determine the SLOP value of the power regulator and set it to rise to about 30°C.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:58 Tuesday 02 November 2021 (18754) Print this report

[Ikeda, Yasui-san, Hirata-san, Ushiba-san].
In the morning we installed the ITMY ribbon heater.
This one does not have the power controller installed yet.

In the afternoon, we installed the PR3 ribbon heater.
We also checked the WIT of the PR3 OpLev to make sure it had not changed significantly before and after the work.

Tests were conducted for PR3, BS, SR2, and SRM.
All tests were started with SLOPE of 4.5 and target temperature of 40°C.
PR3: It was confirmed that the temperature rose to 37°C after 1 hour.
BS: It was confirmed that the temperature rose to 32.5°C after 1 hour.
SR2: It was confirmed that the temperature rose to 27.3 after 1 hour.
SR3: It was confirmed that the temperature rose to 27.9°C after 1 hour.
Since the temperature can be increased by more than +2°C from the room temperature of 23.0-24°C, the requirement is satisfied for the time being.
Thereafter, SLOPE will be adjusted if the required temperature is higher.

[Remaining work on temperature control equipment]
EY: Installation of ribbon heater, installation of thermocouple, wrapping aluminum tape, installation of power regulator, temperature rise test
IY: Installation of power controller, temperature rise test
 => One of the power controller is under repair, so IY was left un-installed for now.
IX, EX, PRM, PR2, PR3, BS, SRM, SR2, SR3: Available
 

Images attached to this comment
Non-image files attached to this comment
takafumi.ushiba - 18:10 Tuesday 02 November 2021 (18755) Print this report

During the work, we keep PR3 aligned and GRX was on.
Attached figure shows the temperature trend and TMS-X_GR_PDA1_OUTPUT.
PD output seems slightly decrease after increasing the temperature PR3, and PD output seems back after recovery of the PR3 temperature.

Note:
There are some delay to change the PD output, especially when decreasing the temperature.
So, it is difficult to confirm that these two are really related with each other.

Images attached to this comment
satoru.ikeda - 15:03 Wednesday 10 November 2021 (18830) Print this report

[Ikeda, Yasui-san]

We installed the ribbon heater and power controller on the ETMY.
As we could not work in the afternoon, we will continue the work later.
 

Non-image files attached to this comment
satoru.ikeda - 20:08 Thursday 11 November 2021 (18848) Print this report

[Ikeda, Yasui-san]

We installed the ETMY power controller and connected the ribbon heater for testing.
We started the test with SLOPE set to 4 and target temperature set to 40°C.
We confirmed that the temperature rise to 34.1°C after about 4 hours.
It was confirmed that the system was operating normally.

We changed the target temperature to 0°C and finished the work.

[Remaining work on temperature controller]
IY: Installation of power controller temperature rise test
 => Power regulator is under repair.
IX, EX, EY, PRM, PR2, PR3, BS, SRM, SR2, SR3: Available

[Main work].
Installation of power regulator: Waiting for power controller to be obtained.
Temperature rise test: Determine the SLOP value of the power controller and set it to rise to about 30°C.
 

Images attached to this comment
Non-image files attached to this comment
Search Help
×

Warning

×